OpenStudy (anonymous):

a man rode a bicycle a straight distance at a speed of 10 miles per hour. he came back the same way, travelling the same distance at a speed of 20 miles per hour. what was the man's total number of miles for the trip back and forth if his total travelling time was 1 hour?

OpenStudy (anonymous):

Let d=the distance from point A to point B. In other words, it's the distance he travels in one direction. Recall that time is equal to distance over velocity. So his time for the forward trip is t=d/10 And the time for his second trip is t=d/20 We know the sum of these times is 1 hour. So d/10 + d/20 = 1. Multiply both sides by 20: 2d+d=20 3d=20 d=20/3 miles Since he goes forward and backward his total distance is 2d or 40/3 miles

OpenStudy (anonymous):

it says here that ur answer is wrong :(

OpenStudy (whpalmer4):

It says here that your spelling of "your" is wrong :-) 20 mph * 1/3 hour = 20/3 miles 10 mph * 2/3 hour = 20/3 miles 20/3 + 20/3 = 40/3 = 13.333 miles The answer is correct.

OpenStudy (dan815):